40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Oct 9 01:29:26 2017, gautam, Update, LSC, DRMI Nosie Budget v3.0 7x
    Reply  Mon Oct 9 11:55:01 2017, Koji, Update, LSC, DRMI Nosie Budget v3.0 
       Reply  Mon Oct 9 22:18:34 2017, gautam, Update, LSC, AS55Q Dark Noise AS55Q_dark.pdf
          Reply  Tue Oct 10 22:04:06 2017, rana, Update, LSC, AS55Q Dark Noise 
          Reply  Wed Oct 11 14:42:03 2017, gautam, Update, LSC, AS55Q Dark Noise AS55Q_darkNoises.pdf
             Reply  Wed Oct 11 19:31:32 2017, gautam, Update, LSC, AS55Q Dark Noise AS55Q_darkNoises.pdf
                Reply  Thu Oct 12 01:50:11 2017, gautam, Update, LSC, AS55Q Dark Noise AS55_dark.png
                   Reply  Thu Oct 12 12:17:28 2017, gautam, Update, LSC, AS55Q Dark Noise D040179_AS55_40m.pdf
                      Reply  Fri Oct 13 12:26:12 2017, gautam, Update, LSC, AS55Q Dark Noise AS55_DemodNoises.pdf
                         Reply  Mon Oct 16 16:01:04 2017, gautam, Update, LSC, AS55Q Dark Noise DemodBoardwOscope.pdfDemodBoardwOscope_ASD.pdf
                            Reply  Tue Oct 17 19:31:53 2017, gautam, Update, LSC, AS55Q Dark Noise D040179_TFs.pdfAS55_DemodNoises.pdf
    Reply  Tue Nov 7 17:45:05 2017, gautam, Update, LSC, DRMI Nosie Budget v3.1 SRCL_MICH_whitenoise_tf.pdfPRCL_MICH_whitenoise_tf.pdfMICH_aux.pdfC1NB_disp_40m_MICH_NB_2017-10-08.pdf
       Reply  Tue Nov 7 22:56:21 2017, gautam, Update, LSC, DRMI locking recovered DRMI_7Nov20178.png
          Reply  Wed Nov 8 00:28:16 2017, gautam, Update, LSC, Laser intensity coupling measurement attempt PSL_RIN.pdf
       Reply  Wed Nov 8 09:37:45 2017, rana, Update, LSC, DRMI Nosie Budget v3.1 
          Reply  Wed Nov 8 09:59:12 2017, gautam, Update, LSC, DRMI Nosie Budget v3.1 
             Reply  Wed Nov 15 01:37:07 2017, gautam, Update, LSC, DRMI low freq. nosie improved C1NB_disp_40m_MICH_NB_2017-11-15.pdfOL_A2L_couplings.pdf
                Reply  Thu Nov 16 00:53:26 2017, gautam, Update, LSC, DRMI noise sub-budgets C1NB_a2l_40m_MICH_NB_2017-11-15.pdfC1NB_aux_40m_MICH_NB_2017-11-15.pdf
Message ID: 13431     Entry time: Thu Nov 16 00:53:26 2017     In reply to: 13428
Author: gautam 
Type: Update 
Category: LSC 
Subject: DRMI noise sub-budgets 

I've incorporated the functionality to generate sub-budgets for the various grouped traces in the NBs (e.g. the "A2L" trace is really the quadrature sum of the A2L coupling from 6 different angular servos).

For now, I'm only doing this for the A2L coupling, and the AUX length loop coupling groups. But I've set up the machinery in such a way that doing so for more groups is easy.

Here are the sub-budget plots for last night's lock - for the OL plot, there are only 3 lines (instead of 6) because I group the PIT and YAW contributions in the function that pulls the data from the nds server, and don't ever store these data series individually. This should be rectified, because part of the point of making these sub-budgets is to see if there is a particularly bad offender in a given group.

I'll do a quick OL loop noise budget for the ITM loops tomorrow.

I also wonder if it is necessary to measure the Oplev A2L coupling from lock to lock? This coupling will be dependant on the spot position on the optic, and though I run the dither alignment servos to minimize REFL_DC, AS_DC, I don't have any intuition for how the offset from center of optic varies from lock to lock, and if this is at all significant. I've been using a number from a measurement made in May. Need to do some algebra...

Attachment 1: C1NB_a2l_40m_MICH_NB_2017-11-15.pdf  412 kB  | Show | Show all
Attachment 2: C1NB_aux_40m_MICH_NB_2017-11-15.pdf  305 kB  | Show | Show all
ELOG V3.1.3-